The sum of two consecutive integers is 77. Find the numbers.

thank you

Answer by stanbon(75887)

The sum of two consecutive integers is 77. Find the numbers.
-----------
1st: x
2nd: x+1
-----------------
Equation:
x + (x+1) = 77
2x = 76
x = 38 (1st number)
x+1 = 39 (2nd number)
=======================
Cheers,
Stan H.